# Duplex-GS: Proxy-Guided Weighted Blending for Real-Time Order-Independent Gaussian Splatting

### v1

@@ VII-A1 Datasets
We conducted extensive experiments on 31 realistic scenes drawn from various datasets, following the configurations in 3DGS [7] and Scafolld-GS [16]. Specifically, we evaluated 13 publicly released datasets, including nine from Mip-NeRF360 [36], two from Tanks and Temples [37] and two from DeepBlending [38]. Additionally, we tested 10 further scenes—eight from BungeeNeRF [40] and two from VR-NeRF [1]—to assess the robustness of our approach in relatively large-scale environments. Scenes from BungeeNeRF feature multi-scale outdoor city scenes for validating view-adaptive rendering, while those from VR-NeRF provide intricate indoor scenes to evaluate detailed rendering capabilities.
@@ VII-A2 Metrics
We evaluate the reconstruction fidelity with PSNR, SSIM [33] and LPIPS [41], which are commonly used in image quality measurement. We also record the model size, frame per second (FPS) as indicators of storage overhead and rendering speed. Additionally, the consumptions of the radix sort are listed to address the effectiveness of our method in reducing the sorting overhead, which is a bottleneck for edge devices.
@@ VII-A3 Baselines
We compare our method with several SOTA approaches, including 3DGS, Mip-NeRF360, Scaffold-GS, Octree-GS, and LC-WSR. Specifically, LC-WSR serves as the baseline for OIT based Gaussian Splatting rendering, while the remaining methods employ the $\alpha$-blending paradigm used in the vanilla 3DGS.
@@ VII-A4 Implementation Details
For the baseline models, we utilize the sum of the $\ell_{1}$ and SSIM losses as the loss function, with the weight $\lambda_{1}$ in Eq. 18 is set to $0.2$. For comparison, the perceptual loss is computed every 20 training iterations, with $\lambda_{2}$ set to $0.5$. For neural Gaussian based models (Scaffold-GS, Octree-GS, and the proposed Duplex-GS), experiments are conducted with encoding dimensions $K$ set to $5$ and $10$ respectively.
For small-scale scenes containing hundreds of training images, training is performed for 40,000 iterations, with densification ending after 25,000 iterations. For large-scale scenes containing thousands of training images, training extends to 100,000 iterations, and densification concludes at 50,000 iterations. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
@@ VII-B Results Analysis
We first validate our methods by showing that it achieves the highest accuracy across several datasets. The results are shown in Tab. I and Tab. II.
@@ VII-B1 Quantitative Results
Tab. I presents a comprehensive comparison of our method with several SOTA baselines on three real-world datasets. On the Mip-NeRF dataset, our approach achieves the highest PSNR when $K=10$, outperforming all baselines. While Mip-Splatting and Octree-GS attain slightly higher SSIM values (0.816), our method yields competitive SSIM (0.813) and achieves improved perceptual quality, as evidenced by a lower LPIPS (0.216). Remarkably, our model also demonstrates strong storage efficiency, requiring significantly less memory than methods such as 3DGS and Mip-Splatting, both of which demand over 750 MB.
On the Tanks&Temples and Deep Blending datasets, although the PSNR values of our approach are marginally lower, it delivers superior visual quality, attaining the highest SSIM and markedly lower LPIPS compared to other models. Notably, when compared to LC-WSR—the OIT variant of Gaussian Splatting—our method achieves better performance across all metrics and datasets, showing the effectiveness of the OIT paradigm in Gaussian Splatting on CUDA based high-end GPUs.
To further assess our method under challenging scenarios involving multi-scale training views and large-scale datasets—common in real-world applications—we conduct experiments on the BungeeNeRF and VR-NeRF datasets (results shown in Table II). On BungeeNeRF, our approach achieves the highest PSNR and lowest LPIPS when $K=10$, and remains highly competitive with $K=5$, attaining the smallest model size among all baselines. The two scenes within VR-NeRF are large-scale, intricate indoor environments containing thousands of training images. Experimental results indicate that OIT based rendering consistently outperforms sort based $\alpha$-blending approaches for such cases, and our proposed method achieves the highest overall accuracy among all evaluated baselines.
@@ VII-B2 Qualitative Results
We present qualitative results and comparisons with SOTA baselines to demonstrate the effectiveness of our proposed method in artifact removal and detail preservation. As illustrated in Fig. 6 using examples from the Mip-NeRF360 dataset, our method eliminates the “transparency” artifacts observed in the previous OIT based Gaussian model, LC-WSR, while maintaining intricate scene details. For multi-scale and large-scale scenarios, as shown in Fig. 7, the proposed Duplex-GS successfully captures fine details, even in textureless and marginal regions. Additionally, our approach effectively removes “popping” artifacts, benefiting from the OIT blending kernel (see Fig. 8).
@@
TABLE III: Efficiency analysis of Duplex-GS and competing baselines.
@@
Ours 184 147 232 124
@@ VII-B3 Efficiency Analysis
Tab. III provides a comparison of rendering speeds, measured in FPS, between our method and several state-of-the-art baselines across four benchmark datasets. Our approach consistently achieves the highest or highly competitive FPS on all datasets. On the Mip-NeRF360 dataset, our method achieves 184 FPS, surpassing all other baselines and comparable to the fastest prior approach, Octree-GS (183 FPS). For the Tanks & Temples dataset, our method obtains 147 FPS, demonstrating performance close to the best-performing baseline. While on the Deep Blending dataset, our FPS of 232 is similarly on par with the leading method. Most notably, on the BungeeNeRF dataset, our method achieves 124 FPS, a considerable improvement over existing baselines. These results highlight the efficiency and scalability of our approach, demonstrating real-time rendering capability and practical superiority in intricate and multi-scale scenarios.
@@
Radix sort is employed in 3DGS to determine the explicit ordering of Gaussian primitives, sorting $b$-digit numbers with $m$ possible values in $\mathcal{O}\left(b\cdot(n+m)\right)$ time and $\mathcal{O}(n+m)$ space. Notably, the sequence length $n$ contributes linearly to the overall complexity. To evaluate the effectiveness of our approach in reducing the computational and memory requirements of view-adaptive sorting, we record the sorting length and corresponding memory consumption during the radix sort procedure. The results, presented in Tab. IV, show that our method successfully mitigates the sorting overhead by introducing cell proxies, thereby demonstrating superior efficiency compared to existing approaches.
@@
TABLE IV: Comparisons about the overhead of radix sort stage. The proposed cell rasterization consistently reduce both the time and space consumptions compared with the vanilla 3DGS rasterization.
@@ VII-C Ablation Studies
VII-C1 Early Termination
The introduction of the early termination mechanism is a key advantage of our method over previous approaches such as LC-WSR. To visualize the mechanism, we record the termination ratio with
@@ VII-C1 Early Termination
where $N^{\text{rendered}}$ denotes the number of Gaussians contributed to the final color, $N^{\text{val}}$ denotes that of the total interacted valid Gaussians along the ray. The results are visualized in Fig. 9. Our method demonstrates termination behavior which closely matches that of the vanilla 3DGS, whereas prior sort-free Gaussian Splatting methods lack this feature. To evaluate its effectiveness, we conduct an ablation study by disabling early termination during model training. As shown in Tab. V, enabling early termination leads to improvements in both accuracy and efficiency.
@@
TABLE V: Ablation experiments.
@@
w/o early termination 27.66 0.801 0.220 155
w/o cell restriction 27.54 0.799 0.223 229
w/o $v_{i}$ 27.59 0.801 0.218 174
@@
VII-C2 Cell Geometry Restriction
Compared with anchor, cell has specific geometry, restricted by Eq. 12 and Eq. 13, to align the regions between cell and the corresponding Gaussians. To show the effects of these conditions, we mute these restriction and the results are shown in Tab. V. The accuracy decreases as expected, however, the rendering speed is improved, which is regarded as the compromise between accuracy and efficiency.
VII-C3 Transmittance Design
As the sorting results of cell level cannot reflect the exact order of Gaussians. we introduce learnable parameter $v_{i}$ to compensate the ambiguity. We train models without $v_{i}$ to show its effectiveness. The results are shown in Tab. V. By introducing $v_{i}$, the accuracy is improved and the efficiency is slightly improved due to the different amount of cell primitives.
@@ VIII Limitation and Conclusion
In this work, we propose a dual-layered Gaussian model that effectively resolves the “popping” artifacts present in vanilla 3DGS and the “transparency” artifacts associated with OIT Gaussian Splatting. Leveraging cell search rasterization and an OIT weighted blending technique, our approach substantially reduces the overhead of global sorting. Implemented with CUDA, our method reconstructs intricate scene details while maintaining real-time rendering speed, thereby demonstrating the practicality and effectiveness of the OIT rendering paradigm for Gaussian Splatting models.
Nevertheless, several limitations remain. Specifically, the geometric relationships between cells and Gaussians are not yet fully explored, leading to performance degradation due to intervals as illustrated in Fig. 5. Additionally, while our current implementation is based on CUDA devices, it does not yet exploit intra-cell parallelism for Gaussian blending. Future work will aim to address these limitations, including efficient implementation on specialized hardware such as Field-Programmable Gate Array (FPGA) and Application-Specific Integrated Circuit (ASIC).

### v2

@@ VII-A1 Datasets
We conducted extensive experiments on scenes from multiple public datasets, adopting the standard configurations established by 3DGS [11].
Our evaluation encompasses 13 standard benchmarks: nine scenes from Mip-NeRF360 [59], two from Tanks&Temples [60], and two from DeepBlending [61]. To further assess robustness, we tested on 10 challenging scenes from BungeeNeRF [63], featuring multi-scale outdoor environments, and VR-NeRF [1], containing large-scale intricate interiors. Additionally, we included the large-scale aerial scene Block_All scene from MatrixCity [64], which comprises 5,621 training and 741 testing images covering an urban area of $2.7~\text{km}^{2}$.
All experiments were performed at a resolution of 1K.
@@ VII-A2 Metrics
We evaluate the reconstruction fidelity with PSNR, SSIM [56] and LPIPS [65], which are commonly used in image quality measurement. We also record the model size, frame per second (FPS) and runtime breakdown as indicators of storage overhead and rendering speed tested on RTX 4090 GPU. Additionally, the consumptions of the radix sort are listed to address the effectiveness of our method in reducing the sorting overhead comparing with $\alpha$-blending, which is a bottleneck for edge devices. We also report FPS on a laptop with RTX 3060 GPU to verify the applicability of our method in mid-range devices.
@@ VII-A3 Baselines
We compare our method with several SOTA approaches, including 3DGS, Mip-NeRF360, Scaffold-GS, Octree-GS, LC-WSR, and StocSplats. Among these, LC-WSR and StocSplats serve as representatives of the OIT based rendering paradigm for Gaussian splatting, while the remaining methods employ the conventional $\alpha$-blending pipeline used in vanilla 3DGS.
The results for StocSplats are sourced directly from its original publication, which reports performance on the Mip-NeRF 360 dataset using an RTX 4090 GPU. As its training code and pre-trained models are not publicly available, we are unable to evaluate its performance on other benchmarks; the corresponding entries in our comparisons are therefore left blank.
@@ VII-A4 Implementation Details
All baseline models are optimized using a combined loss function comprising the $\ell_{1}$ and SSIM losses. The weighting factor $\lambda_{1}$ in Eq. 18 is set to 0.2. In contrast, a perceptual loss is computed every 20 training iterations with a weight $\lambda_{2}$ of 0.5. For neural Gaussian-based methods, including Scaffold-GS, Octree-GS, and our proposed Duplex-GS, experiments are conducted with encoding dimensions $K$ set to 5 and 10, respectively.
Small-scale scenes are trained for 40,000 iterations, with densification concluding at iteration 25,000. Large-scale scenes are trained for 100,000 iterations, with densification ending at iteration 50,000.
@@ VII-B Results Analysis
We first validate our method by showing that it achieves competitive accuracy across all datasets while maintaining high efficiency. The results are shown in Tab. I and Tab. II.
@@ VII-B1 Quantitative Results
Tab. I presents a comprehensive comparison of our method with several SOTA baselines on three real-world datasets. On the Mip-NeRF360 dataset, our approach achieves the highest PSNR when $K=10$, outperforming all baselines. While Mip-Splatting and Octree-GS attain slightly higher SSIM values (0.816), our method yields competitive SSIM (0.813) and achieves improved perceptual quality, as evidenced by a lower LPIPS (0.216). Remarkably, our model also demonstrates strong storage efficiency, requiring significantly less memory than methods such as 3DGS and Mip-Splatting, both of which demand over 750 MB.
On the Tanks&Temples and DeepBlending datasets, although the PSNR values of our approach are marginally lower, it delivers superior visual quality, attaining the highest SSIM or lowest LPIPS compared to other models. Notably, when compared to LC-WSR, the OIT variant of Gaussian Splatting, our method achieves better performance across all metrics and datasets, showing the effectiveness of the OIT paradigm in Gaussian Splatting on CUDA based high-end GPUs.
To further assess our method under challenging scenarios involving multi-scale training views and large-scale datasets, we conduct experiments on the BungeeNeRF, VR-NeRF and MatrixCity datasets (results shown in Tab. II and Tab. III). On BungeeNeRF, our approach achieves the highest PSNR and lowest LPIPS when $K=10$, and remains highly competitive with $K=5$, attaining the smallest model size among all baselines. The two scenes within VR-NeRF are large-scale, intricate indoor environments containing thousands of training images. Experimental results indicate that OIT based rendering consistently outperforms sort-based $\alpha$-blending approaches for such cases. On the MatrixCity dataset, our method achieves the highest accuracy across all metrics (PSNR/SSIM/LPIPS) while maintaining competitive efficiency with the most efficient baseline (128 FPS vs. 119 FPS) and the smallest training time.
@@ VII-B2 Qualitative Results
We present qualitative results and comparisons with SOTA baselines to demonstrate the effectiveness of our proposed method in artifact removal and detail preservation. As illustrated in Fig. 7 and Fig. 12 using examples from the Mip-NeRF360 dataset, our method eliminates the transparency artifacts observed in the previous OIT based LC-WSR, while maintaining intricate scene details. For multi-scale and large-scale scenarios, as shown in Fig. 8, the proposed Duplex-GS successfully captures fine details, even in texture-less and marginal regions. Additionally, our approach effectively removes popping artifacts, benefiting from the OIT blending kernel (see Fig. 11).
@@
TABLE IV: Efficiency analysis of Duplex-GS and competing baselines, reporting the FPS at 1K resolution on RTX 4090 GPU.
@@
StocSplats3 [52] 125 - - -
Ours $(K=5)$ 184 147 232 124
TABLE V: Runtime Breakdown (ms) on MipNeRF360.
Method Stage bicycle bonsai counter flowers garden kitchen room stump treehill AVE
LC-WSR rasterization 7.0177 2.7192 3.2817 3.8094 7.0134 4.5389 2.8045 4.1043 5.7666 4.5618
blending 10.5629 4.2739 4.0905 4.9504 8.9972 6.0351 4.7305 5.2737 8.6236 6.5042
sum 17.5806 6.9932 8.3722 8.7599 16.0106 10.5741 7.5350 9.3779 14.3903 11.0660
Octree-GS decode 0.0027 0.0024 0.0030 0.0035 0.0038 0.0035 0.0018 0.0019 0.0021 0.0027
rasterization 0.4232 0.3942 0.4723 0.3498 0.4280 0.4598 0.4497 0.3083 0.3823 0.4075
sorting 0.6793 0.5985 0.8814 0.4757 0.6933 0.8888 0.7727 0.3011 0.4941 0.6427
blending 1.3343 1.2240 1.4663 1.0070 1.4140 1.5561 1.5236 0.8118 1.0878 1.2695
sum 2.4394 2.2191 2.8231 1.8367 2.5391 2.9082 2.7477 1.4230 1.9663 2.3225
Ours decode 0.0028 0.0027 0.0028 0.0025 0.0032 0.0030 0.0023 0.0023 0.0023 0.0027
rasterization 0.2562 0.2214 0.2387 0.2047 0.2200 0.2377 0.1944 0.2284 0.2580 0.2289
sorting 0.3799 0.4169 0.6360 0.1576 0.2714 0.5857 0.3576 0.1465 0.1925 0.3493
blending 1.7981 1.3834 1.8776 0.9951 1.6763 1.6698 2.1289 0.8722 1.3549 1.5285
sum 2.4371 2.0243 2.7552 1.3600 2.1709 2.4963 2.6832 1.2493 1.8078 2.1093
@@ VII-B3 Efficiency Analysis
Tab. IV provides a comparison of rendering speeds, measured in FPS, between our method and several state-of-the-art baselines across four benchmark datasets. Our approach consistently achieves the highest or highly competitive FPS on all datasets. On the Mip-NeRF360 dataset, our method achieves 184 FPS, surpassing all other baselines and comparable to the fastest prior approach, Octree-GS (183 FPS). For the Tanks&Temples dataset, our method obtains 147 FPS, demonstrating performance close to the best-performing baseline. While on the DeepBlending dataset, our FPS of 232 is similarly on par with the leading method. Most notably, on the BungeeNeRF dataset, our method achieves 124 FPS, a considerable improvement over existing baselines. These results highlight the efficiency and scalability of our approach, demonstrating real-time rendering capability and practical superiority in intricate and multi-scale scenarios.
Tab. V presents a runtime breakdown to elucidate the mechanisms of our hybrid renderer. The analysis reveals that while LC-WSR benefits from hardware graphics pipeline compatibility, it performs suboptimally on TBR architectures due to redundant Gaussian primitives introduced in rendering (exemplified in Fig. 5) and a lack of physical constraints. In contrast, our method achieves a significant acceleration in the rasterization and sorting stages. This gain is directly attributable to our coarse-grained processing of cell proxies, which drastically reduces the number of primitives handled compared to per-Gaussian methods. This efficiency introduces a trade-off: the blending phase exhibits increased latency. This is an expected consequence of our design, which retains sequential blending for cells and may process redundant Gaussians by operating on cell-based groups (of size $K$) rather than individual Gaussians. Despite this localized overhead, our approach achieves the highest overall rendering speed, demonstrating the effectiveness of the proposed hybrid paradigm.
@@
Radix sort is employed in 3DGS to determine the explicit ordering of Gaussian primitives, sorting $b$-digit numbers with $m$ possible values in $\mathcal{O}\left(b\cdot(n+m)\right)$ time and $\mathcal{O}(n+m)$ space. Notably, the sequence length $n$ contributes linearly to the overall complexity. To evaluate the effectiveness of our approach in reducing the computational and memory requirements of view-adaptive sorting, we record the sorting length and corresponding memory consumption during the radix sort procedure. Compared with methods adhering to conventional $\alpha$-blending paradigm, the results, presented in Tab. VI, show that our method successfully mitigates the sorting overhead by introducing cell proxies, thereby demonstrating superior efficiency.
VII-B4 Hyperparameter Analysis
We give analysis on the influence on the hyperparameter $K$, which controls the number of Gaussians decoded per cell. Since the opacity $\sigma$, decoded via $F_{\sigma}$ and activated by a Tanh function, must satisfy $\sigma>0$ for a Gaussian to be considered valid and included in rendering, it directly regulates the number of active primitives. The results show that while varying $K$ leads to only minor deviations in reconstruction accuracy, it substantially impacts computational efficiency. As illustrated in Fig. 9, increasing $K$ results in more Gaussians being activated during rendering, which slows down the blending stage (Fig. 9-a). Conversely, a higher $K$ reduces the number of proxies involved, thereby accelerating the rasterization process (Fig. 9-b). This trade-off elucidates the performance characteristics of our hybrid renderer: the slower blending speed compared to $\alpha$-blending-based frameworks and significant acceleration over LC-WSR are attributable to the number of valid Gaussians (Fig. 9-a), Furthermore, the results confirm that our cell-based framework exploits spatial sparsity more effectively than anchor-based approaches, as evidenced by the fewer proxies located within the same viewing frustum (Fig. 9-b).
Figure 9: Performance analysis with different hyperparameter $K$ on garden scenario from Mip-NeRF 360 dataset, which calculates metric (PSNR [dB]/FPS), the number of valid Gaussians and proxies inside a same viewing frustum during training. (a) The valid Gaussians is number of Gaussians with $\sigma>0$, indicating the efficiency of blending phase. (b) The involved proxies is the number of cells or anchors located inside the viewing frustum. The proposed cell is more sparse than anchors, resulting in acceleration in rasterization stage.
@@ VII-B4 Hyperparameter Analysis
TABLE VI: Comparisons about the overhead of radix sort stage. The proposed cell rasterization consistently reduce both the time and space consumptions compared with $\alpha$-blending based methods.
@@ VII-C Ablation Studies
Figure 10: Visualization of ablation study. Quantitative results are shown in Tab. VII.
VII-C1 Cell vs. Anchor
The introduction of the ellipsoidal cell represents a key innovation, enabling efficient proxy rasterization and physical grounded WSR-based blending. In contrast, anchors adhere to the conventional $\alpha$-blending pipeline, which processes Gaussians individually. To validate the effectiveness of cells in proxy rasterization, we conduct an ablation study by replacing cells with anchor points while maintaining our hybrid rendering paradigm and retraining the models. As shown in Tab. VII, this anchor-based rasterization fails due to the lack of well-defined visible regions, confirming the necessity of our geometric cell structure.
VII-C2 Cell-level Transmittance
The physically grounded transmittance, shown in Eq. 16, is the key innovation which eliminates transparency artifacts as well as enable early termination. We replace it with the conventional LC-WSR blending kernel in the proposed pipeline, which is marked as “w/o $T^{\text{cell}}$” in Tab. VII and Fig. 10. The absence of physical constraints result in transparency artifacts and degraded accuracy.
VII-C3 Early Termination (ET)
The introduction of the early termination mechanism is a key advantage of our method over LC-WSR, which selectively omit computations that have negligible impact on the final rendered output with extremely small transmittance. To visualize the mechanism, we record the termination ratio with
@@ VII-C3 Early Termination (ET)
where $N^{\text{rendered}}$ denotes the number of Gaussians contributed to the final color, $N^{\text{val}}$ denotes that of the total interacted valid Gaussians along the ray. Since the proposed transmittance achieves a decay pattern similar to $\alpha$-blending (Fig. 5), we maintain the same threshold value as used in vanilla 3DGS. The results are visualized in Fig. 13. Our method demonstrates termination behavior which closely matches that of $\alpha$-blending, whereas prior sort-free Gaussian Splatting methods lack this feature. To evaluate its effectiveness, we conduct an ablation study by disabling early termination during model training. As shown in Tab. VII, enabling early termination leads to improvements in both accuracy and efficiency.
@@
TABLE VII: Ablation experiments. Visual examples are shown in Fig. 10.
@@
w/o cell 19.30 0.655 0.365 270
w/o $T^{\text{cell}}$ 26.60 0.790 0.231 212
w/o ET 27.66 0.801 0.220 155
w/o alignment 27.54 0.799 0.223 229
w/o $v_{n}$ 27.59 0.801 0.218 174
@@
VII-C4 Alignment Restriction
Unlike Gaussians whose geometry is directly optimized through blending operations, cells are updated guided by Gaussians and several constraints (Eq. 12 and Eq. 13) are imposed to align coverage between cell and the corresponding Gaussians. To show the effectiveness of these conditions, we mute these restriction and the results are shown in Tab. VII. The accuracy decreases as expected, however, the rendering speed is improved, which offers a solution for resource-constrained platform for its significant acceleration while maintaining competitive visual quality.
VII-C5 Order Confusion
As the sorting results of cell level cannot reflect the exact order of Gaussians. we introduce learnable parameter $v_{n}$ to compensate the ambiguity. We train models without $v_{n}$ to show its effectiveness. The results are shown in Tab. VII. By introducing $v_{n}$, the accuracy is improved as more degree-of-freedom is introduced to handle the overlap ambiguity among cells.
TABLE VIII: Efficiency analysis of Duplex-GS and competing baselines on Mip-NeRF360 datasets, reporting FPS at 1K resolution on a laptop with RTX 3060 GPU .
Method 3DGS LC-WSR Octree-GS Ours (KK=5) Ours (w/o alignment)
FPS 42 21 47 49 72
VII-D Performance on Mid-range GPUs
We test rendering performance on a laptop with RTX 3060 GPU to validate the practicability of our method on mid-range devices. The results are shown in Tab. IV. Our method consistently achieves highest rendering speed while maintaining competitive visual quality.
@@ VIII Limitation and Conclusion
This work presents a dual-hierarchy Gaussian model that effectively resolves popping artifacts inherent in vanilla $\alpha$-blending and the transparency artifacts of WSR. By introducing geometric proxy and a hybrid rendering paradigm, our approach substantially reduces the overhead associated with global sorting while achieving real-time performance as well as competitive visual quality.
Several limitations suggest directions for future work. First, the geometric relationships between cells and their constituent Gaussians require further optimization, as imperfect spatial alignment can lead to performance degradation and redundant calculation, as illustrated in Fig. 6. Second, the core WSR operation within each cell is inherently order-independent, presenting substantial, yet unexplored, parallel computing opportunities. Custom hardware accelerators based on FPGA or ASIC can capitalize on this parallelism to achieve higher performance and energy efficiency compared to general-purpose GPUs, thereby facilitating deployment on resource-constrained edge devices.